OpenStudy (anonymous):

A professor must randomly select 4 students to participate in a mock debate. There are 15 students in his class. In how many different ways can these students be selected, if the order of selection does not matter?

jimthompson5910 (jim_thompson5910):

Let's say you had 4 slots: A,B,C,D how many choices do you have for slot A?

OpenStudy (anonymous):

15

jimthompson5910 (jim_thompson5910):

how many do you have for slot B?

OpenStudy (anonymous):

15

jimthompson5910 (jim_thompson5910):

once you pick a person, you cannot pick them again so you have 15-1 = 14 choices for slot B

jimthompson5910 (jim_thompson5910):

how about slot C?

OpenStudy (anonymous):

13

jimthompson5910 (jim_thompson5910):

and slot D?

OpenStudy (anonymous):

omg i get it now thank you

jimthompson5910 (jim_thompson5910):

what result do you get

OpenStudy (anonymous):

32760

jimthompson5910 (jim_thompson5910):

that would be the answer if order mattered but order does NOT matter

jimthompson5910 (jim_thompson5910):

if you had 4 students, call them A,B,C,D then ABCD is the same as ACBD and ADCB there are 4! = 4*3*2*1 = 24 ways to rearrange ABCD. So we have to divide by 24 to make the correction

jimthompson5910 (jim_thompson5910):

32760/24 = 1,365

jimthompson5910 (jim_thompson5910):

there are 1,365 different ways to pick 4 people (order doesn't matter)

OpenStudy (anonymous):

oh okay thank you

jimthompson5910 (jim_thompson5910):

you're welcome